# DFM-VLA: Iterative Action Refinement for Robot Manipulation via Discrete Flow Matching

### 27 Mar 2026 v1

@@ 1 Introduction
Most existing discrete VLA systems fall into two families. Autoregressive (AR) methods (brohan2022rt; kim2024openvla; univla) decode tokens sequentially with a next-token prediction objective. However, their inherently left-to-right decoding order makes it difficult to revise erroneous tokens once they are emitted. Discrete diffusion (DD) methods (song2025accelerating; liang2025discrete; wen2025llada) improve parallelism and often reduce latency, but they may still produce erroneous tokens in early iterations under confidence-guided decoding (see Figure˜1). As a result, AR and DD paradigms struggle with a shared issue in robot control: early decoding errors cannot be effectively corrected and therefore propagate through the action chunk, degrading downstream robotic task performance. Recent Discrete Flow Matching (DFM) (wang2025fudoki; luo2025next; havasi2025edit; nguyen2025oneflow; deng2025uniform) studies on Large-Language Models (LLMs) or VLMs have demonstrated competitive performance on text reasoning and editing, as well as image and video generation, by enabling iterative token refinement and flexible generation trajectories guided by velocity field. These advances inspire us to transfer the same refinement principle to discrete action generation for robotic control.
@@ 2 Related Works
Discrete diffusion VLA models extend dLLM-style (yu2025discrete) parallel denoising to action sequences, enabling full-attention, multi-token decoding with iterative refinement. Representative methods differ in their corruption processes and efficiency strategies. PD-VLA (song2025accelerating) follows a BART-style noising scheme by substituting tokens with vocabulary items and learning to reconstruct the sequence, while Discrete Diffusion VLA (liang2025discrete) and LLADA-VLA (wen2025llada) adopt BERT-style masking with a dedicated mask token (bert). To reduce inference cost, CEED-VLA (song2025ceed) applies consistency distillation to shrink the number of denoising steps with minimal performance loss. On the data and modeling side, Dream-VLA (ye2025dream) scales discrete diffusion pretraining on the OXE (o2024open) datasets following OpenVLA (kim2024openvla). Likewise, UD-VLA (chen2025unified), dVLA (wen2025dvla), and MM-ACT (liang2025mm) incorporate visual or textual chain-of-thought (CoT)and jointly diffuse future frames, reasoning traces, and action tokens for unified perception–reasoning–action generation. However, these dVLA methods cannot perform token-level refinement, as their updates are restricted to masked positions rather than allowing full-sequence refinement.
@@ 4 Method
This section is organized into four parts. We first introduce the model architecture in Sec. 4.1. We then present two velocity-field constructions: 1) an auxiliary velocity head formulation in Sec. 4.2 and 2) an embedding-guided formulation in Sec. 4.3. Finally, we describe the two-stage inference procedure in Sec. 4.4.
@@ 5.5 In-Depth Analysis
We keep the architecture unchanged and compare AR, DD, and our method, while also validating the effectiveness of Adaptive Cache in inference. Results in Table˜6 show that DFM-VLA achieves the best action quality among the compared decoding strategies. With Adaptive Cache enabled, both DD and DFM obtain clear speedups, and DFM+Adaptive Cache delivers the fastest decoding while maintaining strong performance. These results indicate that DFM provides a favorable quality–efficiency trade-off under matched decoding steps.

### 31 Mar 2026 v2

@@ 1 Introduction
Most existing discrete VLA systems fall into two families. Autoregressive (AR) methods (brohan2022rt; kim2024openvla; univla) decode tokens sequentially with a next-token prediction objective. However, their inherently left-to-right decoding order makes it difficult to revise erroneous tokens once they are emitted. Discrete diffusion ($D$) methods (song2025accelerating; liang2025discrete; wen2025llada) improve parallelism and often reduce latency, but they may still produce erroneous tokens in early iterations under confidence-guided decoding (see Figure˜1). As a result, AR and $D$ paradigms struggle with a shared issue in robot control: early decoding errors cannot be effectively corrected and therefore propagate through the action chunk, degrading downstream robotic task performance. Recent Discrete Flow Matching (DFM) (wang2025fudoki; luo2025next; havasi2025edit; nguyen2025oneflow; deng2025uniform) studies on Large-Language Models (LLMs) or VLMs have demonstrated competitive performance on text reasoning and editing, as well as image and video generation, by enabling iterative token refinement and flexible generation trajectories guided by velocity field. These advances inspire us to transfer the same refinement principle to discrete action generation for robotic control.
@@ 2 Related Works
Discrete diffusion VLA models extend dLLM-style (yu2025discrete) parallel denoising to action sequences, enabling full-attention, multi-token decoding with iterative refinement. Representative methods differ in their corruption processes and efficiency strategies. PD-VLA (song2025accelerating) follows a BART-style noising scheme by substituting tokens with vocabulary items and learning to reconstruct the sequence, while Discrete Diffusion VLA (liang2025discrete) and LLADA-VLA (wen2025llada) adopt BERT-style masking with a dedicated mask token (bert). To reduce inference cost, CEED-VLA (song2025ceed) applies consistency distillation to shrink the number of denoising steps with minimal performance loss. On the data and modeling side, Dream-VLA (ye2025dream) scales discrete diffusion pretraining on the OXE (o2024open) datasets following OpenVLA (kim2024openvla). Likewise, UD-VLA (chen2025unified), dVLA (wen2025dvla), and $M$-ACT (liang2025mm) incorporate visual or textual chain-of-thought (CoT)and jointly diffuse future frames, reasoning traces, and action tokens for unified perception–reasoning–action generation. However, these dVLA methods cannot perform token-level refinement, as their updates are restricted to masked positions rather than allowing full-sequence refinement.
@@ 4 Method
This section is organized into four parts. We first introduce the model architecture in Sec. 4.1. We then present two velocity-field constructions: 1) an auxiliary velocity head formulation in Sec. 4.2 and 2) an embedding-guided formulation in Sec. 4.3. Finally, we describe the two-stage inference procedure in Sec. $4.$
@@ 5.5 In-Depth Analysis
We keep the architecture unchanged and compare AR, $D$, and our method, while also validating the effectiveness of Adaptive Cache in inference. Results in Table˜6 show that DFM-VLA achieves the best action quality among the compared decoding strategies. With Adaptive Cache enabled, both $D$ and DFM obtain clear speedups, and DFM+Adaptive Cache delivers the fastest decoding while maintaining strong performance. These results indicate that DFM provides a favorable quality–efficiency trade-off under matched decoding steps.